2.2 A manufacturer has been ordered by the City of Hoboken to stop discharging waste liquids into the city sewer system. Youc) Calculate the EUAC for the System C. [2 points) d) Which plan should be adopted? Why? (2 point]
Answer #1

a) EUAC of system A = 300,000(A/P, 12%, 30) + 42,000 - 24,000(A/F, 12%, 30)

= 300,000(0.1241) + 42,000 - 24,000(0.00414)

= 37,230 + 42,000 - 99.36

= $79,130.64

b) EUAC of system B = 325,000(A/P, 12%, 30) + 35,000 - 26,000(A/F, 12%, 30)

= 325,000(0.1241) + 35,000 - 26,000(0.00414)

= 40,332.5 + 35,000 - 107.64

= $75,224.86

c) EUAC of system C = 400,000(A/P, 12%, 30) + 28,000 - 32,000(A/F, 12%, 30)

= 400,000(0.1241) + 28,000 - 32,000(0.00414)

= 49,640 + 28,000 - 132.48

= $77,507.52

d) Plan B should be adopted. Because the EUAC of plan B is less than plan A and C.